# Corks, exotic 4-manifolds and genus functions

### v3

@@ 1.4. Exotically knotted submanifolds
Note that manifolds are compact due to our convention, and thus any $n=4$ example is assumed to have non-empty boundary. This question is open even if ambient 4-manifolds are restricted to be simply connected and closed. A partial answer to the $n=4$ case is Corollary 1.13, but, to the best of the author’s knowledge, the corresponding result is not known in the $n=2$ case.
As is well-known, every closed 3-manifold admits an embeddding into a simply connected closed 4-manifold (see e.g. [2, Theorem 2.1]). Hence, for a given closed 3-manifold, it would be natural to ask what is the minimum second Betti number of a simply connected closed 4-manifold for which the 3-manifold is realized as a pair of exotically knotted (or exotically embedded) submanifolds. We show that every non-negative even integer is realized as such a minimum second Betti number by using Theorem 1.14 and a result of Kawauchi [43].
Corollary 1.18.
For each non-negative even integer $n$, there exists a closed $3$-manifold $M$ and a simply connected closed $4$-manifold $Z$ with $b_{2}=n$ such that $Z$ admits infinitely many pairwise exotically knotted and exotically embedded submanifolds diffeomorphic to $M$ and that $M$ admits no embedding into any simply connected closed $4$-manifold with $b_{2}<n$.
We remark that the above $Z$ can be chosen to be $\frac{n}{2}(S^{2}\times S^{2})$ for each $n$. It is likely that our construction gives infinitely many distinct homology cobordism classes of 3-manifolds for each $n$, but we do not pursue this point.
@@ 2. Genus functions
In this section, we introduce a few notions of equivalences for genus functions. We also observe that genus functions can determine smooth structures among symplectic 4-manifolds.
@@
For a second homology class $\alpha$ of $X$, the integer $g_{X}(\alpha)$ is defined to be the minimal genus of a closed surface smoothly embedded in $X$ that represents $\alpha$. The map $g_{X}:H_{2}(X)\to\mathbb{Z}$ is called the genus function of $X$.
This function is well-defined, because every second homology class of $X$ is represented by a closed surface (see [38, Exercise 4.5.12(b)]). We will say that a continuous map $f:X\to Y$ preserves the genus functions if the induced homomorphism $f_{*}:H_{2}(X)\to H_{2}(Y)$ preserves the genus functions, that is, $f_{*}$ satisfies $g_{X}(\alpha)=g_{Y}(f_{*}(\alpha))$ for any $\alpha\in H_{2}(X)$. We here introduce a notion of equivalence for genus functions of 4-manifolds.
@@
The genus functions of $X$ and $Y$ will be called equivalent if there exists a homeomorphism $f:X\to Y$ that preserves the genus functions. If $f$ furthermore preserves the orientations of $X$ and $Y$, then the genus functions will be called orientedly equivalent.
It is easy to see that if two 4-manifolds are diffeomorphic (resp. orientedly diffeomorphic), then their genus functions are equivalent (resp. orientedly equivalent). Hence, the genus function is an invariant of smooth structures.
@@
(1) The genus functions of $X$ and $Y$ will be called algebraically equivalent if there exists an isomorphism $H_{2}(X)\to H_{2}(Y)$ that preserves the genus functions and the intersection forms.
(2) The genus functions of $X$ and $Y$ will be called algebraically equivalent for some orientations if $g_{X}$ is algebraically equivalent to $g_{Y}$ or to $g_{\overline{Y}}$.
(3) For subgroups $F_{X}\subset H_{2}(X)$ and $F_{Y}\subset H_{2}(Y)$, the restrictions of the genus functions of $X$ and $Y$ to $F_{X}$ and $F_{Y}$ will be called algebraically equivalent if there exists an isomorphism $F_{X}\to F_{Y}$ that preserves the genus functions and the intersection forms.
It is clear that if the genus functions of two 4-manifolds are algebraically inequivalent (resp. algebraically inequivalent for any orientations), then they are orientedly inequivalent (resp. (unorientedly) inequivalent) and hence orientedly non-diffeomorphic. We remark that the definition of equivalence in the paper [61] of Stipsicz and Szabó is the same as our algebraic equivalence, not our (ordinary) equivalence.
As we will see, it is useful to consider the genus functions modulo torsion part.
@@
For a class $\widehat{\alpha}$ of $H_{2}(X)/\mathrm{Tor}$, we define an integer $g^{*}_{X}(\widehat{\alpha})$ by
$g^{*}_{X}(\widehat{\alpha})=\min\{g_{X}(\alpha)\mid\textnormal{$\alpha\in H_{2}(X)$ represents $\widehat{\alpha}$.}\}.$
The map $g^{*}_{X}:H_{2}(X)/\mathrm{Tor}\to\mathbb{Z}$ will be called the torsion-free genus function of $X$.
Clearly, $g^{*}_{X}$ is completely determined by $g_{X}$. We note that $g^{*}_{X}$ is often easier to deal with, since adjunction inequalities give lower bounds for values of $g^{*}_{X}$ rather than $g_{X}$.
We define their equivalences and algebraic equivalences in the same way as before.
@@
(1) The torsion-free genus functions of $X$ and $Y$ will be called equivalent if there exists a homeomorphism $f:X\to Y$ such that the isomorphism $H_{2}(X)/\mathrm{Tor}\to H_{2}(Y)/\mathrm{Tor}$ induced by $f$ preserves $g^{*}_{X}$ and $g^{*}_{Y}$. If $f$ furthermore preserves orientations of $X$ and $Y$, then the torsion-free genus functions will be called orientedly equivalent.
(2) The torsion-free genus functions of $X$ and $Y$ will be called algebraically equivalent if there exists an isomorphism $H_{2}(X)/\mathrm{Tor}\to H_{2}(Y)/\mathrm{Tor}$ that preserves $g^{*}_{X},g^{*}_{Y}$ and the intersection forms.
(3) The torsion-free genus functions of $X$ and $Y$ will be called algebraically equivalent for some orientations if $g^{*}_{X}$ is algebraically equivalent to $g^{*}_{Y}$ or to $g^{*}_{\overline{Y}}$.
(4) For subgroups $F_{X}\subset H_{2}(X)/\mathrm{Tor}$ and $F_{Y}\subset H_{2}(Y)/\mathrm{Tor}$, the restrictions of the torsion-free genus functions of $X$ and $Y$ to $F_{X}$ and $F_{Y}$ will be called algebraically equivalent if there exists an isomorphism $F_{X}\to F_{Y}$ that preserves the torsion-free genus functions and the intersection forms.
Let us recall that a rational surface is orientedly diffeomorphic to $\mathbb{C}P^{2}\#n\overline{\mathbb{C}P^{2}}$, and that a ruled surface is diffeomorphic to the total space of an $S^{2}$-bundle over a closed surface or to its blow-ups. We observe that genus functions uniquely determine smooth structures of rational surfaces and ruled surfaces among symplectic 4-manifolds. This fact easily follows from [50, Proposition 4.3].
@@
Let $N_{g}$ denote the total space of a non-trivial $S^{2}$-bundle over the closed surface $\Sigma_{g}$ of genus $g\geq 0$, and let $R$ be a rational surface or a ruled surface. Then $R$ is orientedly diffeomorphic to one of the $4$-manifolds $\mathbb{C}P^{2}\#n\overline{\mathbb{C}P^{2}}$, $(S^{2}\times\Sigma_{g})\#n\overline{\mathbb{C}P^{2}}$ and $N_{g}\#n\overline{\mathbb{C}P^{2}}$ for some non-negative integers $n,g$. We easily see that $R$ contains a closed surface representing a non-torsion second homology class whose self-intersection number is larger than $22g^{\prime}-$ 2, where $g^{\prime}$ denotes the genus of the surface. In the case where $b_{2}^{+}(\overline{R})\geq 1$, the orientation reversal $\overline{R}$ also contains such a surface as seen from its diffeomorphism type.
Now suppose that the genus function of a closed symplectic 4-manifold $X$ is equivalent to $R$. Then we have $b_{2}^{+}(X)\geq 1$, and it thus follows from the above paragraph that $X$ contains a closed surface of genus $g^{\prime}$ representing a non-torsion second homology class whose self-intersection number is larger than $22g^{\prime}-$ 2. Hence, by [50, Proposition 4.3], $X$ is orientedly diffeomorphic to a rational surface or to a ruled surface. We easily see that two 4-manifolds in these classes of 4-manifolds are diffeomorphic if and only if they are homeomorphic, by comparing their homology groups and intersection forms. Since $R$ is also diffeomorphic to one of these 4-manifolds and is homeomorphic to $X$, it follows that $X$ is diffeomorphic to $R$. ∎
@@
3. Corks, homology cobordisms and exotic 4-manifolds
In this section, we review (generalized) corks, homology cobordisms and constructions of exotic 4-manifolds that utilize corks and invertible homology cobordisms. We also give a minor improvement of the method of Akbulut and Ruberman [8] that modifies a generalized cork into exotic 4-manifolds.
@@ 3.1. Generalized corks
(1) Let $C$ be a (not necessarily contractible) 4-manifold and let $f:\partial C\to\partial C$ be a self-diffeomorphism of the boundary $\partial C$ that preserves the orientation. The pair $(C,f)$ will be called a generalized cork if $f$ extends to a self-homeomorphism of $C$, but does not extend to any self-diffeomorphism of $C$. The order of a generalized cork $(C,f)$ is defined to be the minimum positive integer $k$ for which the power $f^{k}$ extends to a self-diffeomorphism of $C$. If $f^{k}$ does not extend to any self-diffeomorphism of $C$ for any positive integer $k$, then $(C,f)$ will be called an infinite order generalized cork. A generalized cork $(C,f)$ is called a cork if $C$ is contractible.
(2) Let $(C,f)$ be a generalized cork, and assume $C$ is embedded in a 4-manifold $X$. We remove $C$ from $X$ and glue it back along the resulting boundary via the gluing map $f$. This operation will be called a generalized cork twist along $(C,f)$.
(3) Two generalized corks $(C,f)$ and $(C,g)$ will be called equivalent if the self-diffeomorphism $g^{-1}\circ f$ of $\partial C$ extends to a self-diffeomorphism of $C$.
@@
3.2. Homology cobordisms
Next, we recall the definitions of a cobordism and related terminologies and fix conventions. Throughout this paper, we denote the interval $[0,1]$ by $I$. Let $M,N$ be closed (possibly disconnected) 3-manifolds, and let $X$ be a 4-manifold having $M$ as an oriented boundary component. For convenience, we allow $N$ to be empty.
@@ 3.2. Homology cobordisms
(1) A (possibly disconnected) 4-manifold $P$ is called a cobordism from $M$ to $N$ if $\partial P$ is orientedly diffeomorphic to the disjoint union $\overline{M}\sqcup N$, and every connected component of $P$ has a non-empty boundary component that belongs to $\overline{M}$. In this case, $M$ and $N$ will be respectively called a negative boundary and a positive boundary of $P$. We often identify $\partial P$ with the disjoint union $\overline{M}\sqcup N$ by fixing a diffeomorphism. For an orientation preserving self-diffeomorphism $f$ of $M$, let $X\cup_{f}P$ denote the 4-manifold constructed from $X$ by attaching $P$ along $M$, where we glue each $x\in M\subset\partial X$ with $f(x)\in M\subset\overline{\partial P}$. When we do not specify the gluing map $f$, we often denote a glued 4-manifold by $X\cup_{M}P$. Unless otherwise stated, we always attach a cobordism to a 4-manifold along a negative boundary.
(2) For a cobordism $P$ from $M$ to $N$, the cobordism $\overline{P}$ from ${N}$ to ${M}$ is called the upside down cobordism of $P$.
@@
(4) A cobordism $P$ from $M$ to $N$ is called invertible if there exists a cobordism $Q$ from $N$ to $M$ such that a glued 4-manifold $P\cup_{N}Q$ is orientedly diffeomorphic to the product $I\times M$. The cobordism $Q$ will be called an inverse cobordism of $P$.

### v5

@@ 1.4. Exotically knotted submanifolds
This question is open even if ambient 4-manifolds are restricted to be simply connected and closed. Corollary 1.13 gives a partial answer to the $n=4$ case, but to the best of the author’s knowledge, it is not known whether the corresponding result holds in the $n=2$ case.
@@ 2. Genus functions
In this section, we introduce notions of equivalences for genus functions, as well as a notion of the torsion-free genus function. We then prove that the values of the genus function do depend on the torsion part of the second homology group, and hence the genus function is not necessarily determined by the torsion-free genus function. Finally, we show that genus functions can determine smooth structures among symplectic 4-manifolds.
@@
For a second homology class $\alpha$ of $X$, the integer $g_{X}(\alpha)$ is defined to be the minimum genus of a closed surface smoothly embedded in $X$ that represents $\alpha$. The map $g_{X}:H_{2}(X)\to\mathbb{Z}$ is called the genus function of $X$.
This function is well-defined, because every second homology class of $X$ is represented by a closed surface (see [39, Exercise 4.5.12(b)]). We occasionally treat disconnected 4-manifolds as well, and for such 4-manifolds, we allow a representative surface to be disconnected and define the genus of such a surface as the sum of the genera of its connected components. Then, we define the genus function of a disconnected 4-manifold in the same way as above. Note that, for a (connected) 4-manifold, these two definitions give the same function as easily seen.
We will say that a continuous map $f:X\to Y$ preserves the genus functions if the induced homomorphism $f_{*}:H_{2}(X)\to H_{2}(Y)$ preserves the genus functions, that is, $f_{*}$ satisfies $g_{X}(\alpha)=g_{Y}(f_{*}(\alpha))$ for any $\alpha\in H_{2}(X)$. We here introduce a notion of equivalence for genus functions of 4-manifolds.
@@
The genus functions of $X$ and $Y$ are called equivalent if there exists a homeomorphism $f:X\to Y$ that preserves the genus functions. If $f$ furthermore preserves the orientations of $X$ and $Y$, then the genus functions are called orientedly equivalent.
It is easy to see that if two 4-manifolds are diffeomorphic (resp. orientedly diffeomorphic), then their genus functions are equivalent (resp. orientedly equivalent). Hence, the genus function is a diffeomorphism invariant of 4-manifolds.
@@
(1) The genus functions of $X$ and $Y$ are called algebraically equivalent if there exists an isomorphism $H_{2}(X)\to H_{2}(Y)$ that preserves the genus functions and the intersection forms.
(2) The genus functions of $X$ and $Y$ are called algebraically equivalent for some orientations if $g_{X}$ is algebraically equivalent to $g_{Y}$ or to $g_{\overline{Y}}$.
(3) For subgroups $F_{X}\subset H_{2}(X)$ and $F_{Y}\subset H_{2}(Y)$, the restrictions of the genus functions of $X$ and $Y$ respectively to $F_{X}$ and $F_{Y}$ are called algebraically equivalent if there exists an isomorphism $F_{X}\to F_{Y}$ that preserves the restrictions of the genus functions and the intersection forms.
It is clear that if the genus functions of two 4-manifolds are algebraically inequivalent, then they are orientedly inequivalent. Hence, if they are algebraically inequivalent for any orientations, then they are inequivalent. We remark that the definition of equivalence in the paper [63] of Stipsicz and Szabó is the same as our algebraic equivalence, not our (ordinary) equivalence.
As we will show in Sections 7 and 8, it is useful to consider the genus function modulo torsion part. We thus introduce the following function.
@@
For a class $A$ of $H_{2}(X)/\mathrm{Tor}$, we define an integer $g^{*}_{X}(A)$ by
$g^{*}_{X}(A)=\min\{g_{X}(\alpha)\mid\textnormal{$\alpha\in H_{2}(X)$ represents $A$.}\}.$
The map $g^{*}_{X}:H_{2}(X)/\mathrm{Tor}\to\mathbb{Z}$ is called the torsion-free genus function of $X$.
Clearly, the torsion-free genus function is completely determined by the genus function. The torsion-free genus function is often easier to deal with, since the adjunction inequality for the Seiberg–Witten invariant gives a lower bound for $g_{X}(\alpha)$ up to torsion. By contrast, here we show that the values of the genus functions do depend on torsion parts even for closed 4-manifolds having non-vanishing Seiberg–Witten invariants, which is of independent interest. Hence, the genus functions are not necessarily determined by the torsion-free genus functions.
Theorem 2.5.
There exists a closed $4$-manifold $X$ with $b_{2}^{+}>1$ having non-vanishing Seiberg-Witten invariant that admits a class $\alpha$ and a torsion class $\tau$ of $H_{2}(X)$ satisfying $g_{X}(\alpha)\neq g_{X}(\alpha+\tau)$. Furthermore, there exist such closed $4$-manifolds having arbitrarily large $b_{2}^{+}$ and $b_{2}^{-}$.
Proof.
Let $Y$ be an arbitrary closed spin 4-manifold, and let $R$ be the closed spin 4-manifold that is the total space of an $S^{2}$-bundle over $\mathbb{R}P^{2}$ (see [39, Exercises 5.7.7.(a)]). We note that $R$ is a rational homology 4-sphere satisfying $H_{2}(R)\cong\mathbb{Z}/2\mathbb{Z}$. We consider the closed 4-manifold $X=Y\#2\overline{\mathbb{C}P^{2}}\#R$. In the case where $Y$ satisfies $b_{2}^{+}>1$ and has a non-vanishing Seiberg–Witten invariant, the 4-manifold $X$ also satisfies these conditions due to the blow-up formula [47, Proposition 2]. Since each elliptic surface $E(2n)$ ($n\in\mathbb{N}$) is an example of such a 4-manifold $Y$, the 4-manifold $X$ realizes arbitrarily large $b_{2}^{+}$ and $b_{2}^{-}$.
Let $e_{1},e_{2}$ be a standard orthogonal basis of $H_{2}(2\overline{\mathbb{C}P^{2}})$. Then, the mod 2 reduction of the class $\alpha=3e_{1}+e_{2}$ of $H_{2}(X)$ is the Poincaré dual of $w_{2}(X)\in H^{2}(X;\mathbb{Z}/2\mathbb{Z})$. Due to Rochlin’s signature theorem, the signature $\sigma(X)$ of $X$ satisfies $\alpha\cdot\alpha-\sigma(X)\equiv-8\not\equiv 0\pmod{16}$. We thus obtain $g_{X}(\alpha)>0$ by [39, Exercises 5.7.7.(b)]). (In fact, $g_{X}(\alpha)=1$ as seen from the rest of this proof.) Let $V$ be the 4-manifold obtained from the 4-ball by attaching a 2-handle along the (10)(-10)-framed figure-eight knot. This manifold V{V} admits an embedding into $2\overline{\mathbb{C}P^{2}}$ that sends a generator of $H_{2}(V)$ to the class $\alpha$ due to a result of Aceto et.al. [1]. Furthermore, by a result of Levine [53, Theorem 1.1 and Remark 1.2], the figure-eight knot bounds a disk in the rational homology 4-ball $R-\mathrm{int}\,D^{4}$. The above inequality thus implies $g_{X}(\alpha+\tau)=g_{2\overline{\mathbb{C}P^{2}}\#R}(\alpha+\tau)=0<g_{X}(\alpha)$, where $\tau$ denotes the torsion class of $H_{2}(R)$. Therefore, the desired claim follows. ∎
We define equivalences for torsion-free genus functions in the same way as before. We state two of them for accuracy.
@@
(1) The torsion-free genus functions of $X$ and $Y$ are called equivalent if there exists a homeomorphism $f:X\to Y$ such that the isomorphism $H_{2}(X)/\mathrm{Tor}\to H_{2}(Y)/\mathrm{Tor}$ induced by $f$ preserves the torsion-free genus functions.
(2) The torsion-free genus functions of $X$ and $Y$ are called algebraically equivalent if there exists an isomorphism $H_{2}(X)/\mathrm{Tor}\to H_{2}(Y)/\mathrm{Tor}$ that preserves the torsion-free genus functions and the intersection forms.
Let us recall that a rational surface is orientedly diffeomorphic to $\mathbb{C}P^{2}\#n\overline{\mathbb{C}P^{2}}$, and that a ruled surface is orientedly diffeomorphic to the total space of an $S^{2}$-bundle over a closed surface or to its blow-ups. We observe that genus functions uniquely determine smooth structures of rational surfaces and ruled surfaces among symplectic 4-manifolds. This fact easily follows from [52, Proposition 4.3].
@@
Let $N_{g}$ denote the total space of a non-trivial $S^{2}$-bundle over the closed surface $\Sigma_{g}$ of genus $g\geq 0$, and let $R$ be a rational surface or a ruled surface. Then $R$ is orientedly diffeomorphic to one of the $4$-manifolds $\mathbb{C}P^{2}\#n\overline{\mathbb{C}P^{2}}$, $(S^{2}\times\Sigma_{g})\#n\overline{\mathbb{C}P^{2}}$ and $N_{g}\#n\overline{\mathbb{C}P^{2}}$ for some non-negative integers $n,g$. Note that $R$ satisfies $b_{2}^{+}=1$. We easily see that $R$ contains a closed surface representing a non-torsion second homology class whose self-intersection number is larger than $2g^{\prime}-2$, where $g^{\prime}$ denotes the genus of the surface. In the case where $b_{2}^{+}(\overline{R})\geq 1$, the orientation reversal $\overline{R}$ also contains such a surface as seen from its diffeomorphism type.
Now suppose that the genus function of a closed symplectic 4-manifold $X$ is equivalent to that of $R$. Then we have $b_{2}^{+}(X)\geq 1$, and it thus follows from the above paragraph that $X$ contains a closed surface of genus $g^{\prime}$ representing a non-torsion second homology class whose self-intersection number is larger than $2g^{\prime}-2$. Hence, by [52, Proposition 4.3], $X$ is orientedly diffeomorphic to a rational surface or to a ruled surface. We see that two 4-manifolds in these classes of 4-manifolds are diffeomorphic if and only if they are homeomorphic, by comparing their homology groups and intersection forms. Since $R$ is also diffeomorphic to one of these 4-manifolds and is homeomorphic to $X$, it follows that $X$ is diffeomorphic to $R$. ∎
@@
3. Corks, invertible cobordisms and exotic 4-manifolds
In this section, we review (generalized) corks, invertible cobordisms and constructions of exotic 4-manifolds that utilize corks and invertible homology cobordisms. Furthermore, we improve the method of Akbulut and Ruberman [8] that modifies generalized corks into exotic 4-manifolds.
@@ 3.1. Generalized corks
(1) Let $C$ be a (not necessarily contractible) 4-manifold and let $f:\partial C\to\partial C$ be a self-diffeomorphism of the boundary $\partial C$ that preserves the orientation. The pair $(C,f)$ is called a generalized cork if $f$ extends to a self-homeomorphism of $C$, but does not extend to any self-diffeomorphism of $C$. The order of a generalized cork $(C,f)$ is defined to be the minimum positive integer $k$ for which the power $f^{k}$ extends to a self-diffeomorphism of $C$. If $f^{k}$ does not extend to any self-diffeomorphism of $C$ for any positive integer $k$, then $(C,f)$ is called an infinite order generalized cork. A generalized cork $(C,f)$ is called a cork if $C$ is contractible.
(2) Let $(C,f)$ be a generalized cork, and assume $C$ is embedded in a 4-manifold $X$. We remove $C$ from $X$ and glue it back along the resulting boundary via the gluing map $f$. This operation is called a generalized cork twist along $(C,f)$.
(3) Two generalized corks $(C,f)$ and $(C,g)$ are called equivalent if the self-diffeomorphism $g^{-1}\circ f$ of $\partial C$ extends to a self-diffeomorphism of $C$.
@@
3.2. Invertible cobordisms
We recall the definitions of an invertible cobordism and related terminologies and fix conventions. Throughout this paper, we denote the interval $[0,1]$ by $I$. Let $M,N$ be closed (possibly disconnected) 3-manifolds, and let $X$ be a 4-manifold having $M$ as an oriented boundary component. For convenience, we allow $N$ to be empty.
@@ 3.2. Invertible cobordisms
(1) A possibly disconnected 4-manifold $P$ is called a ((44-dimensional)) cobordism from $M$ to $N$ if $\partial P$ is orientedly diffeomorphic to the disjoint union $\overline{M}\sqcup N$, and every connected component of $P$ has a non-empty boundary component that belongs to $\overline{M}$. In this case, $M$ and $N$ are respectively called a negative boundary and a positive boundary of $P$. We often identify $\partial P$ with the disjoint union $\overline{M}\sqcup N$ by fixing a diffeomorphism. For an orientation-preserving self-diffeomorphism $f$ of $M$, let $X\cup_{f}P$ denote the 4-manifold constructed from $X$ by attaching $P$ along $M$, where we glue each $x\in M\subset\partial X$ with $f(x)\in M\subset\overline{\partial P}$. When we do not specify the gluing map $f$, we often denote a glued 4-manifold by $X\cup_{M}P$. Unless otherwise stated, we always attach a cobordism to a 4-manifold along a negative boundary of the cobordism.
(2) For a cobordism $P$ from $M$ to $N$, the cobordism $\overline{P}$ from N{N} to M{M} is called the upside down cobordism of $P$.
@@
(4) A cobordism $P$ from $M$ to $N$ is called invertible if there exists a cobordism $Q$ from $N$ to $M$ such that a glued 4-manifold $P\cup_{N}Q$ is orientedly diffeomorphic to the product $I\times M$. The cobordism $Q$ is called an inverse cobordism of $P$.
We use the following terminology.
Definition 3.3.
For a positive integer $n$, attach $n$ 1-handles $h_{1}^{1},h_{2}^{1},\dots,h_{n}^{1}$ to a 4-manifold $X$, and then attach $n$ 2-handles to the resulting 4-manifold $X^{\prime}$. The set of these 1- and 2-handles is called homotopically canceling pairs of $1$- and $2$-handles if the link in $\partial X^{\prime}$ consisting of the attaching circles of 2-handles is homotopic to a link consisting of knots $K_{1},K_{2},\dots,K_{n}$ in $\partial X^{\prime}$, such that each $K_{i}$ intersects the belt sphere of the 1-handle $h_{i}^{1}$ geometrically exactly once.
Here we give a simple construction of an invertible cobordism as well as its properties. We will later give much more flexible constructions, see Lemma 6.7.
Lemma 3.4.
Suppose a cobordism $P$ from $M$ to $N$ is obtained from $I\times M$ by attaching homotopically canceling pairs of $1$- and $2$-handles, none of which are attached to $0\times M$. Then $P$ is invertible, and the upside down cobordism $\overline{P}$ is its inverse cobordism. Moreover, $P\cup_{\mathrm{id}_{N}}\overline{P}$ is diffeomorphic to $I\times M$ fixing the boundary $\overline{M}\sqcup M$ pointwise, where $\overline{M}\sqcup M$ is identified with $\{0,1\}\times M$ via the natural identification.
Proof.
We consider the 5-manifold $P\times I$. This manifold is obtained from $(I\times M)\times I$ by attaching homotopically canceling pairs of $1$- and $2$-handles in $1\times M\times I$. Since a homotopy of attaching circles of these 2-handles can be approximated by a smooth isotopy, we can eliminate these pairs. We thus have a diffeomorphism between $P\times I$ and $(I\times M)\times I$ that fixes $(0\times M)\times I$ pointwise. By restricting this diffeomorphism to the exteriors of $(0\times M)\times I$ in the boundaries of these two 5-manifolds, we see that $P\cup_{\mathrm{id}_{N}}\overline{P}$ is diffeomorhic to $I\times M$ fixing the boundary $\overline{M}\sqcup M$ pointwise. ∎
Remark 3.5.
Under the assumption of this lemma, $P$ is a homology cobordism. Furthermore, the homotopy extension property (see [40, Proposition 0.18]) implies that $P$ admits a deformation retraction onto the boundary component $M$.